9) (5z-2)(5z+2)....use foil....need some help here...thanks
Found 2 solutions by stanbon, dini:
Answer by stanbon(75887)

(5z-2)(5z+2)....use foil
-----
F: 5z*5z = 25z^2
O: 5z*2 = 10z
I: 5z*-2 =-10z
L: -2*2 = -4
---------------------------
Combine the four results to get:
25z^2 - 4
===============
Cheers,
Stan H.

Answer by dini(5)

(5z-2)*(5z+2)
=(5z)^2+2(5z)-2(5z)-2*2
=25z^2-4